# Solving Harder Reinforcement Learning Problems

### v2

@@ 1 Introduction
The $\mathrm{pass@k}$ was championed by [11] who gave a popular unbiased estimator of the metric which we derive from a new perspective (to set up our gradient estimators) as Theorem 1, generalize to continuous rewards in Theorem 3, and provide additional characterisations in Corollaries 2 and 3. Concurrently with our work [39] offered an elegant variance reduction method for the gradient of the $\mathrm{pass@k}$ which corresponds to the special case $n=k$ of our Equation 33. Interpreting $\mathrm{pass@k}$ in terms of a partial sort, [14] and [43] present elegant approximations that are rather general but less efficient in our setting. Others have provided variational approximations for handling the closely related Best-of-NN [10, 6] and other more general [7] inference-time algorithms. The contrasting idea of training a model to approximate the Best-of-NN prediction with a single sample was addressed by [34]. Our contribution can be interpreted as a generalization and variance reduction of [39]. For a general discussion of gradient estimation, variance reduction, and Monte Carlo, we recommend [29, 31].
@@ 2.1 An Unbiased $\mathrm{pass@k}$ Estimator
where (6) holds because the sum on the r.h.s. of (5) is the number of subsets of size $k$ of the (nc)(n-c) incorrect elements. Since averaging in this way retains unbiasedness, this completes the proof. ∎ We show in Corollary 2 that no such unbiased estimator exists for $n\leq k$, and in Corollary 3 that the asymptotic variance of this estimator decreases at a rate of $1/n$.
@@ A.3 Proof of Theorem 2
If $f_{i}=1$ then all subsets that include $i$ also include at least one correct element (ii itself), so that $m^{(1)}$ is just the number of subsets of size $k$ of $\{1,2,\dots,n\}$ that include $i$, which equals the number of subsets of size k1k-1 of $\{1,2,\dots,n-1\}$:
@@ Appendix C Additional Figures
all subsets baselined: $s^{(\mathrm{loo})}$ — our novel estimator of Equation 29 that analytically sums over all subsets of size $k$ of the $n$ samples with our unbiased baseline method that subtracts for each element $i$ the mean of the estimator over all subsets of size $k$ that do not include ii.

### v3

@@ 1 Introduction
The $\mathrm{pass@k}$ was championed by [11] who gave a popular unbiased estimator of the metric which we derive from a new perspective (to set up our gradient estimators) as Theorem 1, generalize to continuous rewards in Theorem 3, and provide additional characterisations in Corollaries 2 and 3. Concurrently with our work [39] offered an elegant variance reduction method for the gradient of the $\mathrm{pass@k}$ which corresponds to the special case $n=k$ of our Equation 33. Interpreting $\mathrm{pass@k}$ in terms of a partial sort, [14] and [43] present elegant approximations that are rather general but less efficient in our setting. Others have provided variational approximations for handling the closely related Best-of-$N$ [10, 6] and other more general [7] inference-time algorithms. The contrasting idea of training a model to approximate the Best-of-$N$ prediction with a single sample was addressed by [34]. Our contribution can be interpreted as a generalization and variance reduction of [39]. For a general discussion of gradient estimation, variance reduction, and Monte Carlo, we recommend [29, 31].
@@ 2.1 An Unbiased $\mathrm{pass@k}$ Estimator
where (6) holds because the sum on the r.h.s. of (5) is the number of subsets of size $k$ of the (nc)(n-c) incorrect elements. Since averaging in this way retains unbiasedness, this completes the proof. ∎ We show in Corollary 2 that no such unbiased estimator exists for $n<k$, and in Corollary 3 that the asymptotic variance of this estimator decreases at a rate of $1/n$.
@@ A.3 Proof of Theorem 2
If $f_{i}=1$ then all subsets that include $i$ also include at least one correct element ($i$ itself), so that $m^{(1)}$ is just the number of subsets of size $k$ of $\{1,2,\dots,n\}$ that include $i$, which equals the number of subsets of size k1k-1 of $\{1,2,\dots,n-1\}$:
@@ Appendix C Additional Figures
all subsets baselined: $s^{(\mathrm{loo})}$ — our novel estimator of Equation 29 that analytically sums over all subsets of size $k$ of the $n$ samples with our unbiased baseline method that subtracts for each element $i$ the mean of the estimator over all subsets of size $k$ that do not include $i$.